This is NOT from a text book.
The perimeter of one square is 36 inches more than the perimeter of a second square, and the area of the first square is 243 square inches more than the area of the second square. Find the area of the larger square.
My work so far:
Square #1 Perimeter=4s+36 Area=s^2+243
Square #2 Perimeter= 4s Area=s^2
I am unsure of how to set up the equation.

Answer by scott8148(6628)

sq#1 ___ side length = perimeter / 4 = s+9 ___ area = (side length)^2 = (s+9)^2 = s^2 + 243

s^2 + 18s + 81 = s^2 + 243

18s = 162 ___ s = 9

A = 9^2 + 243 = 324